# STELLAR: Structure-guided LLM Assertion Retrieval and Generation for Formal Verification

### v1

@@ Abstract.
conference: Design Automation Conference; July 26-29, 2026; Long Beach, CAjournalyear: 2026
@@ 1. Introduction
Formal Verification (FV) is vital in the VLSI design flow, addressing the growing complexity of hardware architectures (gupta1992formal). In particular, FV applies mathematical methods to prove that a design satisfies a given property, offering stronger guarantees than simulation. A widely used approach is Assertion-Based Verification (ABV) (foster2008assertion; witharana2022survey), which relies on SystemVerilog Assertions (SVAs) to formally specify the expected behavior of a design. SVAs can be verified either statically with formal tools or dynamically in simulation. The effectiveness of FV depends heavily on the quality of SVAs (witharana2022survey). However, manually writing high-quality SVAs is time-consuming, error-prone, and difficult to scale, as it requires translating often ambiguous natural language specifications (SPEC) into precise formal properties. Automating assertion generation is crucial for fully realizing the potential of FV in modern hardware verification. The main challenge lies in bridging the semantic gap between the high-level design intent, which is often described ambiguously in SPEC, and the low-level implementation details in the Register-Transfer Level (RTL) code. Verification engineers must mentally synthesize information from both sources to create high-quality assertions, a process that is challenging to automate.
@@
Despite the capabilities of large language models (LLMs) in code generation, a recent study indicates that commercial LLMs alone are not yet fully equipped for assertion generation, as they often produce syntactic or semantic errors (pulavarthi2025llms). Other recent work has shown potential improvement for LLM-based automating assertion generation. For example, SANGAM (gupta2025sangam) performs iterative reasoning over multimodal specifications and RTL, while AssertionForge (bai2025assertionforge) constructs structured knowledge from SPEC and RTL to synthesize comprehensive design information. Despite these advancements, both works share a key limitation as they aim to create assertions from scratch by only analyzing the underlying design files (SPEC and RTL), without the benefit of learning from existing expert-written assertions in industry codebases. For instance, AssertionForge provides limited context to the LLM regarding the information it extracts from the current design’s SPEC and RTL, including code snippets and information based on matching signal names. This “invent from scratch” approach overlooks the most valuable asset in the industrial FV setting, which is codebases of expert-written RTL and SVA pairs from past/existing projects. Failing to utilize these proven examples results in several limitations including wasted efforts, the reintroduction of previously resolved bugs, and the failure to enforce the consistent, industry-standard SVA styles found in expert-written codebases.
@@ 2.1. Automated Assertion Generation
The manual effort and expertise required for high-quality assertion generation have motivated the research into automating SVA generation (kande2024security; kang2025fveval; liu2024domain; orenes2021autosva). Non-LLM methods such as GOLDMINE (vasudevan2010goldmine), A-Team (danese2017team), and HARM (germiniani2022harm) rely on data mining, templates, and/or static analysis. While effective in certain domains, they often fail to scale to large industrial designs due to the algorithmic complexity of static analysis (pulavarthi2025llms) and state explosion, which limits their capability on large designs (witharana2022survey).
LLM-based approaches have gained a lot of research attention recently (sun2023towards; maddala2024laag; menon2025openassert; ankireddy2025lasa). ChIRAAG (mali2024chiraag) automates assertion generation from natural language SPECs using GPT-4 with iterative refinement. Likewise, AssertLLM (yan2025assertllm) uses multiple specialized LLMs to extract information from SPECs and generate SVAs, with RTL only used for verification. SANGAM (gupta2025sangam) uses multi-modal SPEC files, including diagrams, with Monte Carlo Tree Search for iterative refinement for SVA generation. SANGAM (gupta2025sangam) uses RTL for signal-name mapping between RTL and SPEC and final verification, whereas AssertionForge (bai2025assertionforge) combines SPEC files and RTL into a unified Knowledge Graph to guide the LLM for SVA generation. [^2] Knowledge Graph is a graph-structured representation of entities (nodes) and relations (edges). A critical limitation of prior LLM-based methods is that they generate SVAs from scratch or use context limited to the target design’s SPECs and RTL.
@@ 2.2. Retrieval-Augmented Generation (RAG)
LLMs are powerful but prone to hallucination when not grounded in a reliable context. RAG (lewis2020retrieval), originating in natural language processing tasks, mitigates hallucination by retrieving relevant context from an external knowledge base so that context generation is guided by prior knowledge rather than the model’s world knowledge. RAG uses embedding models to convert text or code into dense vector representations that capture semantic meaning. At retrieval time, RAG finds the items whose embeddings are closest to the query (e.g., by cosine similarity), providing semantically related context to the LLM even when the wording differs.
@@ 2.3. LLM Interaction Paradigms
Table 1 shows the different modes that LLMs can be used. The zero-shot setting relies solely on pretrained knowledge. RAG, rather than relying solely on pretraining, improves reliability by grounding the model with relevant context retrieved from an external knowledge base. Fine-tuning adapts model weights using task-specific data, which can improve performance but (i) requires large, curated labeled datasets; (ii) is computationally expensive; (iii) may cause catastrophic forgetting when learning new information; and (iv) can reduce cross-task generalizability. In hardware verification, RAG emerged recently. AssertionForge (bai2025assertionforge) retrieves context via semantic similarity from RTL and SPEC. However, as shown in Fig. 1, semantic similarity alone often miss the structural control flow essential for assertions, motivating structure-aware retrieval.
@@ 2.4. Hardware Similarity and Structural Representations
Measuring hardware similarity has been explored at gate level (li2022deepgate) and RTL (yu2021hw2vec; akyash2025simeval). HW2VEC (yu2021hw2vec) converts RTL into graph and uses GNNs to generate embeddings. While effective for classification tasks (e.g., detecting Trojans), it lacks the fine-grained structural resolution needed for SVA retrieval. SimEval (akyash2025simeval) evaluates LLM-generated RTL code using syntax, control-flow, and functionality, but oversimplifies results into a single scalar score. If multiple candidates receive same scores for a given query, SimEval provides no tie-break, rendering distinct designs indistinguishable.
@@ 2.5. Dataset Generation for Assertion Learning
A major challenge in automating assertion generation is the lack of large, standardized datasets, as open-source projects often use inconsistent assertion styles (menon2025vert). Nonetheless, hardware designs are typically verified at the module level, which aligns well with the context limits of LLMs. VERT (menon2025vert) addresses this gap by providing the first large-scale benchmark of 20,000 paired RTL code and SVAs, drawn from open-source System-on-Chip designs (zhao2020sonicboom; lee2016rocketchip; xu2022xiangshan), with augmentation to diversify variable names and control structures. VERT systematically includes patterns such as deeply nested if–else statements, synchronous/asynchronous blocks, where LLMs often fails (menon2025vert). STELLAR uses VERT as the knowledge base, enabling scalable structure-aware retrieval of similar (RTL, SVA) pairs.
@@ 3.1. Structural Fingerprinting
Challenge  1 Semantic vs. Structural Similarity. The primary challenge in retrieving relevant examples for SVA generation is choosing the correct similarity metric. Semantic similarity, common in RAG, depends on textual content and variable names but can be misleading in hardware design due to arbitrary signal names that do not reflect logic. Relying solely on semantics leads to retrieving structurally divergent code that happens to share variable names (as shown in Fig. 1), providing misleading context to the LLM. In comparison, structural similarity is a more reliable indicator. However, not all methodologies are effective (Sec. 2.4). GNN-based models such as HW2VEC (yu2021hw2vec) produce coarse embeddings, resulting in vector collapse and collisions, whereby distinct RTL blocks whose embeddings have perfect cosine similarity are mapped to identical vectors. To quantify this, we performed experiments to analyze the embedding space of HW2VEC on 1K randomly selected designs from the VERT dataset. The result is shown in Fig. 2. The fact that 97.7% of the designs were involved in at least one collision confirms that HW2VEC embeddings are unsuitable for distinguishing logic and accurate retrieval.
Figure 2. Collision comparison for HW2VEC (yu2021hw2vec) embeddings and STELLAR structural fingerprinting (lesser is better).
@@ 3.2. Structural Embeddings and Retrieval
Solution  3. To enable efficient similarity search, the structural fingerprints are converted into dense vector embeddings using the lightweight, efficient all-MiniLM-L6-v2 Sentence Transformer (allminilm-l6-v2), suitable for encoding short, structured texts into a semantic space where proximity indicates similarity.
@@ 4.1. Implementation Details
STELLAR uses PyVerilog (Takamaeda2015Pyverilog) for AST parsing. LLM temperature is set to 0 for deterministic outputs , and a 1024-token limit. Performance of STELLAR is evaluated by incorporating the top-3/top-5 most structurally similar examples into the structure-aware prompt, along with target RTL code.
@@ 4.2. Knowledge Base Curation and Stratification
Our experiments leverage the latest version of VERT dataset (menon2025vert) as the foundational source. A preliminary analysis revealed that a subset of these pairs contained syntactic errors. The following examples show two such cases that were filtered out: one with an extra parenthesis and another with an invalid event control placement within a Boolean expression.
@@ 4.4. Evaluation Metrics
We evaluate generated assertions on a stratified query set, comparing with and without STELLAR across three LLMs using top-3 and top-5 retrieved examples. Performance is measured by three metrics: (i) syntax correctness, (ii) similar to (kang2025fveval), BLEU score for lexical similarity to golden SVA and BERT score for semantic similarity to the reference SVAs [^3] Since VERT provides a canonical reference assertion for each RTL snippet, these similarity metrics serve as lightweight proxies for functional alignment and correctness relative to the intended behavior encoded with respect to that reference., (iii) the execution paths covered by the generated assertions for each RTL block.
@@ 5.1. Retrieval Quality
Since the VERT dataset (menon2025vert) lacks a ground truth for RTL similarity, our retrieval mechanism is evaluated with a controlled synthetic experiment. Specifically, we select 100 random RTL modules from the knowledge base for duplication and run two cases: (1) exact duplication (0% renaming), and (2) identifier names are changed (30% renaming) while preserving functionality. For each case, we query the retrieval system to assess whether it can rank the original source high. The results are shown in Table LABEL:tab:robustness_alpha, with N denoting the top-N retrieved results, Recall@N measuring the success rate of retrieving the correct item within the top N results, MRR@N reporting how early the first correct item appears, and nDCG@N measuring ranking quality. All values are normalized to the range of [0, 1]. The higher, the better.
As shown in Table LABEL:tab:robustness_alpha, when no renaming is applied, the baseline that relies on semantic embedding of raw RTL code perfectly retrieves the matching source. However, its performance degrades significantly (Recall@10 drops from 1.0 to 0.2) when 30% of the variables are renamed. This is because the baseline, which directly encodes raw RTL using all-MiniLM-L6-v2 (allminilm-l6-v2) without fingerprinting, is highly sensitive to the change in code lexical representation. In contrast, STELLAR uses structural fingerprinting invariant to variable names, allowing it to remain stable under both settings. These results also show that STELLAR retrieves and ranks relevant RTL codes high even when substantial naming differences are introduced, highlighting its superior practicality to baseline.
@@ 5.2. STELLAR Assertion Generation Quality
Having seen the effectiveness of STELLAR’s retrieval mechanism, we now evaluate its impact on the SVA generation. We compare the performance of STELLAR (using top-3 and top-5 retrieved examples, denoted by K) against a zero-shot baseline [^4] Although AssertionForge (bai2025assertionforge) is the closest prior work, the necessary materials for comparison were not available at the time of submission., where the LLM generates SVAs on RTL without any context provided.

### v2

@@ Abstract.
conference: Conference name; Date; Locationjournalyear: 2026
@@ 1. Introduction
Formal Verification (FV) is vital in the VLSI design flow, addressing the growing complexity of hardware architectures (Gupta, 1992). In particular, FV applies mathematical methods to prove that a design satisfies a given property, offering stronger guarantees than simulation. A widely used approach is Assertion-Based Verification (ABV) (Foster, 2008; Witharana et al., 2022), which relies on SystemVerilog Assertions (SVAs) to formally specify the expected behavior of a design. SVAs can be verified either statically with formal tools or dynamically in simulation. The effectiveness of FV depends heavily on the quality of SVAs (Witharana et al., 2022). However, manually writing high-quality SVAs is time-consuming, error-prone, and difficult to scale, as it requires translating often ambiguous natural language specifications (SPEC) into precise formal properties. Automating assertion generation is crucial for fully realizing the potential of FV in modern hardware verification. The main challenge lies in bridging the semantic gap between the high-level design intent, which is often described ambiguously in SPEC, and the low-level implementation details in the Register-Transfer Level (RTL) code. Verification engineers must mentally synthesize information from both sources to create high-quality assertions, a process that is challenging to automate.
@@
Despite the capabilities of large language models (LLMs) in code generation, a recent study indicates that commercial LLMs alone are not yet fully equipped for assertion generation, as they often produce syntactic or semantic errors (Pulavarthi et al., 2025). Other recent work has shown potential improvement for LLM-based automating assertion generation. For example, SANGAM (Gupta et al., 2025) performs iterative reasoning over multimodal specifications and RTL, while AssertionForge (Bai et al., 2025) constructs structured knowledge from SPEC and RTL to synthesize comprehensive design information. Despite these advancements, both works share a key limitation as they aim to create assertions from scratch by only analyzing the underlying design files (SPEC and RTL), without the benefit of learning from existing expert-written assertions in industry codebases. For instance, AssertionForge provides limited context to the LLM regarding the information it extracts from the current design’s SPEC and RTL, including code snippets and information based on matching signal names. This “invent from scratch” approach overlooks the most valuable asset in the industrial FV setting, which is codebases of expert-written RTL and SVA pairs from past/existing projects. Failing to utilize these proven examples results in several limitations including wasted efforts, the reintroduction of previously resolved bugs, and the failure to enforce the consistent, industry-standard SVA styles found in expert-written codebases.
@@ 2.1. Automated Assertion Generation
The manual effort and expertise required for high-quality assertion generation have motivated the research into automating SVA generation (Kande et al., 2024; Kang et al., 2025; Liu et al., 2024; Orenes-Vera et al., 2021). Non-LLM methods such as GOLDMINE (Vasudevan et al., 2010), A-Team (Danese et al., 2017), and HARM (Germiniani and Pravadelli, 2022) rely on data mining, templates, and/or static analysis. While effective in certain domains, they often fail to scale to large industrial designs due to the algorithmic complexity of static analysis (Pulavarthi et al., 2025) and state explosion, which limits their capability on large designs (Witharana et al., 2022).
LLM-based approaches have gained a lot of research attention recently (Sun et al., 2023; Maddala et al., 2024; Menon et al., 2025b; Ankireddy et al., 2025). ChIRAAG (Mali et al., 2024) automates assertion generation from natural language SPECs using GPT-4 with iterative refinement. Likewise, AssertLLM (Yan et al., 2025) uses multiple specialized LLMs to extract information from SPECs and generate SVAs, with RTL only used for verification. SANGAM (Gupta et al., 2025) uses multi-modal SPEC files, including diagrams, with Monte Carlo Tree Search for iterative refinement for SVA generation. SANGAM (Gupta et al., 2025) uses RTL for signal-name mapping between RTL and SPEC and final verification, whereas AssertionForge (Bai et al., 2025) combines SPEC files and RTL into a unified Knowledge Graph to guide the LLM for SVA generation. [^2] Knowledge Graph is a graph-structured representation of entities (nodes) and relations (edges). A critical limitation of prior LLM-based methods is that they generate SVAs from scratch or use context limited to the target design’s SPECs and RTL.
@@ 2.2. Retrieval-Augmented Generation (RAG)
LLMs are powerful but prone to hallucination when not grounded in a reliable context. RAG (Lewis et al., 2020), originating in natural language processing tasks, mitigates hallucination by retrieving relevant context from an external knowledge base so that context generation is guided by prior knowledge rather than the model’s world knowledge. RAG uses embedding models to convert text or code into dense vector representations that capture semantic meaning. At retrieval time, RAG finds the items whose embeddings are closest to the query (e.g., by cosine similarity), providing semantically related context to the LLM even when the wording differs.
@@ 2.3. LLM Interaction Paradigms
Table 1 shows the different modes that LLMs can be used. The zero-shot setting relies solely on pretrained knowledge. RAG, rather than relying solely on pretraining, improves reliability by grounding the model with relevant context retrieved from an external knowledge base. Fine-tuning adapts model weights using task-specific data, which can improve performance but (i) requires large, curated labeled datasets; (ii) is computationally expensive; (iii) may cause catastrophic forgetting when learning new information; and (iv) can reduce cross-task generalizability. In hardware verification, RAG emerged recently. AssertionForge (Bai et al., 2025) retrieves context via semantic similarity from RTL and SPEC. However, as shown in Fig. 1, semantic similarity alone often miss the structural control flow essential for assertions, motivating structure-aware retrieval.
@@ 2.4. Hardware Similarity and Structural Representations
Measuring hardware similarity has been explored at gate level (Li et al., 2022) and RTL (Yu et al., 2021; Akyash and Mardani Kamali, 2025). HW2VEC (Yu et al., 2021) converts RTL into graph and uses GNNs to generate embeddings. While effective for classification tasks (e.g., detecting Trojans), it lacks the fine-grained structural resolution needed for SVA retrieval. SimEval (Akyash and Mardani Kamali, 2025) evaluates LLM-generated RTL code using syntax, control-flow, and functionality, but oversimplifies results into a single scalar score. If multiple candidates receive same scores for a given query, SimEval provides no tie-break, rendering distinct designs indistinguishable.
@@ 2.5. Dataset Generation for Assertion Learning
A major challenge in automating assertion generation is the lack of large, standardized datasets, as open-source projects often use inconsistent assertion styles (Menon et al., 2025a). Nonetheless, hardware designs are typically verified at the module level, which aligns well with the context limits of LLMs. VERT (Menon et al., 2025a) addresses this gap by providing the first large-scale benchmark of 20,000 paired RTL code and SVAs, drawn from open-source System-on-Chip designs (Zhao et al., 2020; Lee et al., 2016; Xu et al., 2022), with augmentation to diversify variable names and control structures. VERT systematically includes patterns such as deeply nested if–else statements, synchronous/asynchronous blocks, where LLMs often fails (Menon et al., 2025a). STELLAR uses VERT as the knowledge base, enabling scalable structure-aware retrieval of similar (RTL, SVA) pairs.
@@ 3.1. Structural Fingerprinting
Challenge  1 Semantic vs. Structural Similarity. The primary challenge in retrieving relevant examples for SVA generation is choosing the correct similarity metric. Semantic similarity, common in RAG, depends on textual content and variable names but can be misleading in hardware design due to arbitrary signal names that do not reflect logic. Relying solely on semantics leads to retrieving structurally divergent code that happens to share variable names (as shown in Fig. 1), providing misleading context to the LLM. In comparison, structural similarity is a more reliable indicator. However, not all methodologies are effective (Sec. 2.4). GNN-based models such as HW2VEC (Yu et al., 2021) produce coarse embeddings, resulting in vector collapse and collisions, whereby distinct RTL blocks whose embeddings have perfect cosine similarity are mapped to identical vectors. To quantify this, we performed experiments to analyze the embedding space of HW2VEC on 1K randomly selected designs from the VERT dataset. The result is shown in Fig. 2. The fact that 97.7% of the designs were involved in at least one collision confirms that HW2VEC embeddings are unsuitable for distinguishing logic and accurate retrieval.
Figure 2. Collision comparison for HW2VEC (Yu et al., 2021) embeddings and STELLAR structural fingerprinting (lesser is better).
@@ 3.2. Structural Embeddings and Retrieval
Solution  3. To enable efficient similarity search, the structural fingerprints are converted into dense vector embeddings using the lightweight, efficient all-MiniLM-L6-v2 Sentence Transformer (Sentence-Transformers, 2021), suitable for encoding short, structured texts into a semantic space where proximity indicates similarity.
@@ 4.1. Implementation Details
STELLAR uses PyVerilog (Takamaeda-Yamazaki, 2015) for AST parsing. LLM temperature is set to 0 for deterministic outputs , and a 1024-token limit. Performance of STELLAR is evaluated by incorporating the top-3/top-5 most structurally similar examples into the structure-aware prompt, along with target RTL code.
@@ 4.2. Knowledge Base Curation and Stratification
Our experiments leverage the latest version of VERT dataset (Menon et al., 2025a) as the foundational source. A preliminary analysis revealed that a subset of these pairs contained syntactic errors. The following examples show two such cases that were filtered out: one with an extra parenthesis and another with an invalid event control placement within a Boolean expression.
@@ 4.4. Evaluation Metrics
We evaluate generated assertions on a stratified query set, comparing with and without STELLAR across three LLMs using top-3 and top-5 retrieved examples. Performance is measured by three metrics: (i) syntax correctness, (ii) similar to (Kang et al., 2025), BLEU score for lexical similarity to golden SVA and BERT score for semantic similarity to the reference SVAs [^3] Since VERT provides a canonical reference assertion for each RTL snippet, these similarity metrics serve as lightweight proxies for functional alignment and correctness relative to the intended behavior encoded with respect to that reference., (iii) the execution paths covered by the generated assertions for each RTL block.
@@ 5.1. Retrieval Quality
Since the VERT dataset (Menon et al., 2025a) lacks a ground truth for RTL similarity, our retrieval mechanism is evaluated with a controlled synthetic experiment. Specifically, we select 100 random RTL modules from the knowledge base for duplication and run two cases: (1) exact duplication (0% renaming), and (2) identifier names are changed (30% renaming) while preserving functionality. For each case, we query the retrieval system to assess whether it can rank the original source high. The results are shown in Table LABEL:tab:robustness_alpha, with N denoting the top-N retrieved results, Recall@N measuring the success rate of retrieving the correct item within the top N results, MRR@N reporting how early the first correct item appears, and nDCG@N measuring ranking quality. All values are normalized to the range of [0, 1]. The higher, the better.
As shown in Table LABEL:tab:robustness_alpha, when no renaming is applied, the baseline that relies on semantic embedding of raw RTL code perfectly retrieves the matching source. However, its performance degrades significantly (Recall@10 drops from 1.0 to 0.2) when 30% of the variables are renamed. This is because the baseline, which directly encodes raw RTL using all-MiniLM-L6-v2 (Sentence-Transformers, 2021) without fingerprinting, is highly sensitive to the change in code lexical representation. In contrast, STELLAR uses structural fingerprinting invariant to variable names, allowing it to remain stable under both settings. These results also show that STELLAR retrieves and ranks relevant RTL codes high even when substantial naming differences are introduced, highlighting its superior practicality to baseline.
@@ 5.2. STELLAR Assertion Generation Quality
Having seen the effectiveness of STELLAR’s retrieval mechanism, we now evaluate its impact on the SVA generation. We compare the performance of STELLAR (using top-3 and top-5 retrieved examples, denoted by K) against a zero-shot baseline [^4] Although AssertionForge (Bai et al., 2025) is the closest prior work, the necessary materials for comparison were not available at the time of submission., where the LLM generates SVAs on RTL without any context provided.